Determine whether each of these functions is a bijection from R to R.

f(x)=2x+1

f(x)=x^2+1 


1
Expert's answer
2021-01-26T04:41:00-0500

Here,

  • "f(x)=2x+1"

Let "x_1,x_2 \in\R" and let us assume "f(x_1)=f(x_2)"

So,


"f(x_1)=f(x_2)\\
\Rightarrow2x_1+1=2x_2+1\\
\Rightarrow x _1=x_2"

Hence, we have "f(x_1)=f(x_2)" implies "x_1=x_2" .

So, "f" is one-one (injective).


Also we know


"-\infin<x<\infin\\
\Rightarrow -\infin<2x<\infin\\
\Rightarrow -\infin<2x+1<\infin\\
\Rightarrow -\infin<f(x)<\infin"

So, we clearly observe the Co-Domain is the same as the Range, so "f(x)" is surjective.

And hence "f(x)" is bijective.



  • "f(x)=x^2+1"

Here we have


"-\infin<x<\infin\\
\Rightarrow 0\leq x^2<\infin\\
\Rightarrow 1\leq x^2+1<\infin\\
\Rightarrow 1\leq f(x)<\infin"

So, the Co-Domain of "f" is "\R" , but the range of "f" is "[1,\infin)", so "f(x)" is not surjective.

Hence we conclude that "f(x)" is not bijective.